OpenStudy (anonymous):

set up and simplify an integral that represents the area of the surface obtained by rotating the curve x=sin^2(t), y=sin(3t), 0<=t<=(pi/3) around the x-axis

OpenStudy (anonymous):

\[0 \le t \le \pi/3\]

OpenStudy (anonymous):

first express everything like function of x